OpenStudy (anonymous):

need help algebra 2. 3 ^3(sqrt)(2a)-6 ^3(sqrt)(2a)

OpenStudy (gorv):

3^3(sqrt)(2a)-(3.2)^3(sqrt)(2a)

OpenStudy (anonymous):

\[3\sqrt[3]{2a}-6\sqrt[3]{2a}\]

OpenStudy (gorv):

3^3(sqrt)(2a)(1-2^3(sqrt)(2a))

OpenStudy (anonymous):

that is?

OpenStudy (gorv):

\[3^{\sqrt[3]{2a}}(1-2^{\sqrt[3]{2a}})\]

OpenStudy (anonymous):

i mean is that like the answer or a step because im lost on how to do this

OpenStudy (anonymous):

expand the parenthesis. when multiplying add exponents. \[\huge {\color{black} {3^{\sqrt[3]{2a}}-6^{\sqrt[3]{2a}}} }\] THIS?

OpenStudy (anonymous):

\[\huge {\color{black} {-3^{\sqrt[3]{2a}}} }\] would be the result if you subtract because they are like terms.